8. Determine the displacement vector that must be added to the displacement
(25 ̂ - 16 ̂) m to give a displacement of 7.0 m pointing in the +x-direction?
1
Expert's answer
2020-09-14T10:22:51-0400
"(7,0)=(25,-16)+\vec{d}\\\vec{d}=(7-25,16)=(-18,16)\ m"